which type of conic section is given by the following equation? (y-2)^2/3^2-(x+1)^2/5^2=1
Answer by DSMLMD(16)

(y - 2)^2/3^2 - (x + 1)^2/5^2 = 1
(y - 2)^2/9 - (x + 1)^2/25 = 1


from that equation, we can identify easily because we only see the denominator of the equation and a plus (+) and a minus (-) sign of the equation. If they have denominator in the equation that contain the nominator x and y, then they have 2 possibility equation, between ellipse and hyperbola. But, if the middle two fractions are separate and there is a plus (+) sign, it's ellipse. If the middle two fractions are separate and there is a minus (-) sign, it's hyperbola.


So, from the equation (y - 2)^2/9 - (x + 1)^2/25 = 1 the type of conic section is hyperbola.